# Texture-aware and Shape-guided Transformer for Sequential DeepFake Detection

### v2

@@ 4.2 Results
Table 1 shows the performance of our method compared to other counterparts, including Multi-Cls, DETR Carion et al. (2020), DRN Wang et al. (2019), MA Zhao et al. (2021b), Two-Stream Luo et al. (2021), SeqFake-Former Shao et al. (2022) respectively. Multi-Cls and DETR serve as baseline methods adapted from models for general vision tasks. Multi-Cls performs multi-label classification by directly classifying the manipulated images into multiple classes. DETR is a modified object detector that replaces object queries with manipulation annotation. DRN, MA, and Two-Stream are DeepFake detection methods adapted by substituting the binary classification head with a multi-classification head. These methods are all referrenced from Shao et al. (2022). SeqFake-Former is a recent dedicated method for sequential DeepFake detection. We reproduce this method rigorously using their release codes. As shown in Table 1, our method outperforms baseline methods by a large margin, averaging $71.92\%$ and $54.06\%$ in Fixed-Acc and Adaptive-Acc scores, respectively. Moreover, our method surpasses the DeepFake detection methods in both tracks. Notably, compared to SeqFakeFormer, our method achieves improvements by $3.82\%,5.47\%$ and $1.28\%,1.32\%$ in these two tracks.
The results also indicate that all methods perform better in Fixed-Acc than in Adaptive-Acc. This observation is consistent with the findings in Shao et al. (2022) that predicting manipulation sequences with varying lengths is a more challenging task. In real-world applications, manipulation sequences with varying lengths are more practical, thus the improvement of our method in this metric sufficiently represents the practicability of the proposed modules.
Intriguingly, we observe that ResNet-50 performs slightly worse than ResNet-34 across all methods. This contrasts with the expectation that a larger model should have stronger capacity. We discuss this phenomenon in Sec. 4.4.
4.3 Ablation Study
Our method consists of several key components, including the Diversified Pixel Difference Attention module (DPDA), Bidirectional Interactive Cross-attention (BICA), Inverted Order Prediction (IOP), and Shape-guided Gaussian Mapping (SGM). To evaluate the effectiveness of each component, we conduct a series of ablation studies on the facial-components track using the stem of ResNet-34 and ResNet-50.
Specifically, we designed different models based on the following settings: 1) A baseline model with a Transformer architecture that only consists of basic encoders and decoders. 2) Adding the DPDA module to the baseline model, which generates two spatial features from the encoder. Then we directly add these two features and feed them into the sequential decoder. 3) Adopting the BICA strategy on the baseline model for feature integration. 4) Incorporating inverted order prediction. 5) Inserting the SGM module into BICA strategy.
The results of the ablation studies are shown in Table 2. It can be observed that each component contributes to the overall performance. Taking Resnet-34 as an example. The DPDA module improves the Fixed-Acc and Adaptive-Acc by $0.59\%$ and $1.44\%$, respectively. The BICA strategy further enhances the performance by $0.76\%$ and $0.8\%$ in Fixed-Acc and Adaptive-Acc. After applying IOP, the performance is significantly improved by $2.32\%$ in Fixed-Acc and $3.98\%$ in Adaptive-Acc, showing the effectiveness of this strategy in gradually capturing manipulation traces. Finally, the integration of SGM yields further performance, achieving the best performance when all components are applied.
Figure 5: The visualization of our model and SeqFakeFormer under the (a) eyebrow-eye-hair-nose-lip and (b) nose-eyebrow-lip-eye-hair manipulating annotations.
Table 3: Performance on Different CNN Stems.
Stem Fixed-Acc Adaptive-Acc
SFF Ours SFF Ours
ResNet-18 68.41 69.97 49.55 52.26
ResNet-34 70.88 74.70 52.96 58.43
Resnet-50 69.41 73.32 50.22 56.08
Resnet-101 68.87 73.46 49.06 56.77
4.4 Discussion
Attention Visualization. To assess the learning ability of our method, we visualize the attention maps of our method compared with Seq-FakeFormer under two different manipulation annotations, eyebrow-eye-hair-nose-lip and nose-eyebrow-lip-eye-hair. Fig. 5 shows the Grad-CAM Selvaraju et al. (2019) of these methods. We observe that under the same annotation, our method can identify the manipulated regions more precisely. In contrast, Seq-FakeFormer can localize the major manipulated regions but may overlook small manipulations with subtle traces. This result highlights the effectiveness of our method in capturing sequential manipulation traces.
@@ 4.4 Discussion
Furthermore, we also observe that our method can highlight regions that are not annotated as manipulation. We attribute this to the imperfect nature of GAN in the feature space, causing subtle influences on some neighboring areas during the multi-step facial manipulation process. Our method can capture this subtle alternation as evidence for exposing the sequential manipulations.
Controversy Regarding CNN Stem. From Table 3 we observe an unexpected phenomenon where using ResNet-34 as CNN stem achieves better performance than ResNet-50. To further investigate this, we validate our method using ResNet-18 and ResNet-101. Table 3 shows the performance comparison of our method and SeqFake-Former (SFF) on these architectures. It can be seen that the performance of ResNet-18 is notably degraded compared to ResNet-34, aligning with our expectations. However, when using ResNet-101 as the Stem, it yields lower performance than ResNet-50. This finding contradicts the general understanding that increasing the model size and parameters can improve feature extraction capacity. One possible explanation for this phenomenon is that the length of the manipulation sequence is relatively short (less than five), which increases the likelihood of overfitting in large models. This suggests the need to strike a balance between model size and performance in the future.
@@ 5 Conclusion
We introduce a novel Texture-aware and Shape-guided Transformer to enhance detection performance. Compared to the classic Transformer-based methods, we introduce four improvements. Firstly, we describe a new Diversiform Pixel Difference Attention module (DPDA) to capture subtle manipulation traces. Moreover, we introduce a Bidirectional Interaction Cross-attention module (BICA) to explore deep correlations between spatial and sequential features. To further enhance the cross-attention, we describe a Shape-guided Gaussian mapping strategy to provide the priors of manipulation shapes. Lastly, we improve the performance by inverting the prediction order from forward to backward. The results obtained on public datasets demonstrate the efficacy of our method in sequential DeepFake detection.

### v4

@@ 4.2 Results
Table 1 compares the performance of our method with other counterparts, including Multi-Cls, DETR [2], DRN [29], MA [39], Two-Stream [18], SeqFake-Former [23] respectively. Multi-Cls and DETR serve as baseline methods adapted from models for general vision tasks. Multi-Cls performs multi-label classification by directly classifying the manipulated images into multiple classes. DETR is a modified object detector that replaces object queries with manipulation annotation. DRN, MA, and Two-Stream are DeepFake detection methods adapted by substituting the binary classification head with a multi-classification head. SeqFake-Former [23] is a dedicated method for sequential DeepFake detection. MMNet [31] is the latest work achieving state-of-the-art performance. The results of Multi-Cls, DETR, DRN, MA, and Two-stream are referred from [23]. Note that only DRN, MA, Two-stream, and SeqFake-Former have released their codes. For better evaluation, we reproduce them with their released codes (marked by *) (More details in Supplementary). As shown in Table 1, our method outperforms others in both tracks, averaging $72.67\%$ and $55.44\%$ in FACC and AACC scores on ResNet-50. Compared to SeqFake-Former, our method improves by $2.88\%,4.37\%$ and $0.94\%,1.58\%$ in these two tracks on ResNet-50. Moreover, our method averagely surpasses MMNet by $1.07\%$ and $1.8\%$ in FACC and AACC.
In Challenging Scenarios. In real-world scenarios, images often undergo various post-processing operations. To these scenarios, we perform post-processing operations on the facial-components and facial-attributes datasets. Specifically, for each image, we randomly select one or two processes from a set that includes adding Gaussian noises, shifting RGB channels, compression, converting to gray, and jittering colors. This increases the difficulty of original datasets, which we refer to as Facial-components and Facial-attributes. We then evaluate different methods under these challenging scenarios. As shown in Table 2, while all methods suffer a notable performance drop, our model still outperforms others, demonstrating its robustness against external disturbances.
4.3 Analysis
Effect of Each Component. Our method consists of several key components, including the Diversified Pixel Difference Attention module (DPDA), Multi-source Cross-attention (MSCA), Inverted Order Prediction (IOP), and Shape-guided Gaussian Mapping (SGM). To evaluate the effectiveness of each component, we conduct a series of ablation studies on the facial-components track using the stem of ResNet-34 and ResNet-50. Specifically, 1) the baseline model (BL) contains only basic encoders and decoders. 2) Adding the DPDA module to the baseline model and directly feeding the features from the encoder into the sequential decoder. 3) Employ the MSCA strategy for feature integration. 4) Incorporating inverted order prediction. 5) Inserting the SGM module into the MSCA strategy. The results shown in Table 3 indicate the contribution of each component to the overall performance. For instance, with Resnet-34, the DPDA module improves the FACC and AACC by $0.59\%$ and $1.44\%$, respectively. The MSCA strategy further enhances the performance by $0.76\%$ in FACC and $0.8\%$ in AACC. The IOP significantly boosts FACC by $2.32\%$ and AACC by $3.98\%$, showing its effectiveness in gradually capturing manipulation traces. Finally, the integration of all components leads to the best performance.
Generalizability of IOP. The proposed IOP strategy is applicable to many architectures. To verify this, we evaluate the performance of different methods with/without IOP. As shown in Table 4, all methods have gained enhancement on average. Even the methods of MA, DRN, and two-stream that are not specifically designed for sequence detection tasks surprisingly improve their performance by a notable margin. In particular, by applying IOP, Two-Stream even achieves competitive performance with MMnet on the facial-components track, demonstrating the generic effect of IOP for SDD.
Table 5: Different loss weights $\lambda_{1},\lambda_{2}$.
$\lambda_{1}$ $\lambda_{2}$ FACC AACC
0.1 0.1 72.21 56.02
0.01 0.01 73.32 56.84
0.5 0.01 74.20 58.21
1 0.01 73.71 57.24
0.1 0.01 75.21 59.74
Study on Loss Weights $\lambda_{1},\lambda_{2}$. We study the effect of various loss weight $\lambda_{1}$ and $\lambda_{2}$. The results are shown in Table 5. It can be seen that our method is not very sensitive to the values of loss weights (±$(\pm 3\%)$. This is because the classification loss primarily influences the learning direction during training. Compared to the KLD loss $\lambda_{2}$, the reconstruction loss $\lambda_{1}$ seemly has more impact, as it directly instructs the training of encoder ${\cal E}$ in a self-supervised manner. In the main experiment, we use $\lambda_{1}=0.1,\lambda_{2}=0.01$.
Adaptation on One-step DeepFake Detection (ODD). Since one-step detection does not require sequential clues, existing SDD methods can not be directly applied to this task. For more comprehensive studies, we adapt these methods by replicating the one-step label $N=5$ times as the sequential annotations, where $N$ is the maximum manipulation length in dataset. For example, the real faces are assigned with annotations of $[0,0,0,0,0]$, while forged faces correspond to $[1,1,1,1,1]$. For comparison, we train these methods on the FF++ dataset [21] under this configuration. Note the basenetwork of SeqFake-Former and ours is ResNet-34. The results are shown in Table 6, revealing that our method still achieves the best results.
Table 6: Performance on FF++.
Methods ACC
Two-stream 82.40
DRN 81.42
MA 82.64
SeqFake-Former 82.43
TSOM 83.70
Figure 6: Attention visualization.
@@ 4.3 Analysis
Attention Visualization. Fig. 6 shows the Grad-CAM [22] of different methods. We observe that under the same annotation, our method can identify the manipulated regions more precisely. In contrast, Seq-FakeFormer can localize the major manipulated regions but may overlook small manipulations with subtle traces. This result highlights the effectiveness of our method in capturing sequential manipulation traces. Furthermore, we also observe that our method can highlight regions that are not annotated as manipulation. We attribute this to the imperfect nature of GAN in the feature space, causing subtle influences on some neighboring areas during the multi-step facial manipulation process. Our method can capture this subtle alternation as evidence for exposing the sequential manipulations.
@@ 5 Conclusion
This paper introduces a new Transformer design, called TSOM, for sequential DeepFake detection. Our method is inspired by three perspectives: Texture, Shape, and Order of Manipulations, leading to four key improvements. Firstly, we propose a text-aware branch featuring new Diversiform Pixel Difference Attention modules (DPDA) to capture subtle manipulation traces. Secondly, we introduce a Multi-source Cross-attention module (MSCA) to explore deep correlations between spatial and sequential features. To enhance the cross-attention further, we describe a Shape-guided Gaussian Mapping (SGM) strategy to incorporate priors on manipulation shapes. Lastly, we improve performance using a simple Inverted Order prediction (IOP) strategy, which reverses the manipulation annotation order from forward to backward. The results on public datasets demonstrate the efficacy of our method in sequential DeepFake detection.
Acknowledgement. This work is supported in part by the National Natural Science Foundation of China (No.62402464), Shandong Natural Science Foundation (No.ZR2024QF035), China Postdoctoral Science Foundation (No.2021TQ0314, No.2021M703036).